OpenStudy (anonymous):

The 5th term of an arithmetic series is -18 and the 9th term is 4. Find the sum of the first 20 terms.

OpenStudy (campbell_st):

do you know the formula for a term in an arithmetic sequence..?

OpenStudy (anonymous):

tn=ti+(n-1)d

OpenStudy (campbell_st):

ok... I use a instead of ti.... for the 1st value so you know -18 = a + (5 - 1) d or -18 = a + 4d 4 = a + ( 9 -1) d or 4 = a + 8d you need to solve the equations -18 = a + 4d 4 = a + 8d to find the 1st term and the common difference... then you will be able to find the sum of the 1st 20 terms

OpenStudy (anonymous):

Yup, I did that- and I got -3.5 for the d value, then I plugged that into on of the equations and got 32 for the a value (are these values correct?) and then I'm unsure as to where to go from there...

OpenStudy (campbell_st):

so using elimination you may want to check my calculations... -18 = a + 4d - 4 = a + 8d ---------------- -22 = -4d d = -5.5

OpenStudy (campbell_st):

oops d = 5.5

OpenStudy (campbell_st):

once you know the 1st term and the common difference the sum of an arithmetic series is \[S_{n} = \frac{n}{2}[2a + (n - 1)d]\] just substitute and evaluate

OpenStudy (anonymous):

Ah yes indeed, thank you for that correction! Alright, but how would we substitute into the formula, given the fact that we have two terms?

OpenStudy (anonymous):

Hello?

OpenStudy (campbell_st):

ok so a = -40 and d = 5.5 so \[S_{20} = \frac{20}{2}[2 \times -40 + (20 - 1) \times 5.5]\] now just calculate it out.

OpenStudy (anonymous):

Cool, may I ask what formula that is though? Or did you just derive it for this problem in particular? Because I thought the following equation is the one you use for arithmetic\[Sn=(n/2)(a+tn)\] series:

OpenStudy (campbell_st):

well its the same formula as you use you use n/2( 1st + last)